Consider a beam of light with intensity I0.

Imagine the beam passes through two slits 1.50 μm apart and is shined on a screen 55.0 cm from the slits. If the bright fringes on the screen are 20.0 cm apart, what color is the light?
Answer #1

Given,

\\Distance\ between\ slits,\ d=1.50\mu m=1.50\times 10^{-6}m \\Distance\ from\ slits\ to\ screen,\ D=55cm=0.55m \\Fringe\ width,\ \beta =20cm=0.2m

(Distance between two adjacent bright or dark fringes is called as fringe width)

In interference, the fringe width is given by,

\\\beta =\frac{D}{d}\lambda \\ \\\therefore \lambda =\beta \frac{d}{D} \\ie,\ \lambda =0.2\times \frac{1.50\times 10^{-6}}{0.55}=545.45\times 10^{-9}m=545.45nm

This wavelength of 545nm corresponds to GREEN (Emerald green to be precise)